• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Bosted

    Log entry time 23:56:19 on November 3,2002

    Entry number 89533

    keyword=Shift Summary

    We had smooth running until 19:30. Then we had major DAQ problems and
    in the end had to reboot adaql2. Resumed running at 19:30. At 21:20, MCC called
    to ask if we knew anything about the pressure in the Burn-through
    Monitor near the Beam DUmp, as they thought it might be causing some of
    the frequent beam trips (about one every five to 10 minutes). Eric said we
    didn't know about this, and would look into it tommorrow. Meanwhile,
    we have acceptable beam most of the time. The beam was pretty good for
    the rest of the shift.